On Year 1 ABC Company sell equipment to XYZ Company for $415,000 in an installment sale. ABC Company has a basis in the equipment of $250,000. In year 1 ABC company will received three payments totaling $300,000 and in year 2 the final payment totaling $115,000.

  1. What is the gross profit on the sale?
  2. What is the gross profit percentage?
  3. How much taxable income is recognized in year 1?
  4. How much taxable income is recognized in year 2?
  5. If the equipment was sold for a price of $100,000 rather than $400,000, how much loss would be recognized in year 2?
Answer #1

Gross Profit = Sale price - Basis = $415,000 - $250,000 = $165,000

Gross Profit Percentage = (Gross Profit / Sale Price) * 100 = ($165,000 / $415,000) * 100 = 39.76% (approx.)

Taxable income to be recognized in year 1 = $300,000 * 39.76% = $119,280

Taxable income to be recognized in year 2 = $115,000 * 39.76% = $45,724

Need more details to answer the last question.